OpenStudy (anonymous):

Marthas employee benefits include family health care coverage. She contributes 19% of the cost. Martha gets paid biweekly and $131.00 is taken out of each check for family health care. How much does her employer contribute annually for the family coverage?

OpenStudy (anonymous):

Biweekly... So 2 times the number of weeks in a year times her paycheck is the annual amount she makes.

OpenStudy (anonymous):

Multiply this result by (100%-19%)/(19%)

OpenStudy (anonymous):

Basically:\[ \frac{x}{1-0.19} = \frac{52\times 2\times $131.00}{0.19} \]

OpenStudy (anonymous):

Where \(1-0.19\) is the percentage that the employer pays.

OpenStudy (nincompoop):